OpenStudy (anonymous):

On the graph of the equation 8x-3y=24, what is the value of the y-intercept?

OpenStudy (lgbasallote):

let x = 0 then solve for y so.. 8(0) - 3y = 24 what is the value of y then?

OpenStudy (anonymous):

y=8

OpenStudy (anonymous):

@godorovg do not provide just answers . you need this http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

I was going to type why?

OpenStudy (anonymous):

Thank you. & people do it all the time..

OpenStudy (lgbasallote):

why didnt you say so?? i could've randomly said y = 0 and eased all my troubles

OpenStudy (anonymous):

the reason why y=8 is because 3y=24 you / y=and have 24/3 which is y=8

OpenStudy (anonymous):

Lol, does it equal 8 or 0?

OpenStudy (anonymous):

Never mind 0 isn't an answer choice xD

OpenStudy (lgbasallote):

you sure?

OpenStudy (lgbasallote):

because i could've sworn my random answer was right

OpenStudy (anonymous):

ha ha

OpenStudy (anonymous):

@godorovg you provided y=8 ?? if you are not sure of something then do not reply .

OpenStudy (lgbasallote):

anyway @Deedles i'll let you in a secret...if you just answered y = 8 you're dead wrong...that's what happens when you just listen to answers with no solutions...

OpenStudy (anonymous):

if x=0 and you have 3y=24 to get y my itself you divide 3 by 24 which gives you eight

OpenStudy (lgbasallote):

sure...stand by your answer..

OpenStudy (anonymous):

well if i am not right show me why??

OpenStudy (anonymous):

because unless someone proves otherwise I will stand by the answer..

OpenStudy (anonymous):

hint read the question again guys, I know I am right..

OpenStudy (lgbasallote):

HAHAHAHAHA what an ironic comment

OpenStudy (anonymous):

okay are you guys joking with me??

OpenStudy (anonymous):

ha ha ha funny thanks I need a laugh

OpenStudy (anonymous):

oh but the payback opps

OpenStudy (lgbasallote):

let me say it in your own words... "hint read the question again "

OpenStudy (anonymous):

My exam said your wrong xD

OpenStudy (anonymous):

@godorovg now happy ?

OpenStudy (lgbasallote):

8x - 3y = 24 substitute x with 0 8(0) - 3y = 24 -3y = 24 dividee both sides by -3 y = -8 there's your wanted proof @godorovg

OpenStudy (anonymous):

Oh jeez! Hush!

OpenStudy (anonymous):

@godorovg just tell me what is 30/(-6 )= ????

OpenStudy (anonymous):

okay so i didn't see the - sorry

OpenStudy (anonymous):

y=-8

OpenStudy (anonymous):

okay guys I was wrong okay move on I made a mistake it happes

OpenStudy (anonymous):

@lgbasallote I dont CARE what you think, I asked for help cause I know I CAN'T do it. So gtfo. I didn't ask for your secret.

OpenStudy (anonymous):

Ok :)